Point Q is on line segment \overline{PR}

PR


. Given QR=2x+4,

QR=2x+4, PQ=x,

PQ=x, and PR=4x-10,

PR=4x−10, determine the numerical length of \overline{PR}.

PR


.


1
Expert's answer
2021-09-12T16:44:47-0400

Note that "PR=PQ+QR"

Substitute "PR=4x-10" , "PQ=x" , and "QR=2x+4" into the relation:

"4x-10=x+2x+4"

Add like terms:

"4x-10=3x+4"

Subtract "3x" from both sides:

"4x-10-3x=3x+4-3x"

Combine like terms:

"x-10=4"

Add "10" to both sides:

"x=4+10"

Add the numbers:

"x=14"

Note that "PR=4x-10"

Substitute "x=14" into the equation:

"PR=4(14)-10=46"

The numerical length of "\overline{PR}" is "46"